ABC’s last dividend paid was $7.68, its required return is 21%, its growth rate is 6%. What is ABC's expected stock price in 11 years?
Answer #1

Current price=D1/(Required return-Growth rate)

=(7.68*1.06)/(0.21-0.06)

=$54.272

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence expected price =$54.272*(1.06)^11

=$54.272*1.898298558

=$103.02(Approx).